# Continuous Invariance Learning

### v1

@@ Appendix D Discussions on the first moments
To encourage conditional independence $\boldsymbol{y}\perp\boldsymbol{t}|\Phi(\boldsymbol{x})$, existing IRM variants try to align $\mathbb{E}^{t}[\boldsymbol{y}|\Phi(\boldsymbol{x})]$ (see Section 2.1) and our CIL proposes to align $\mathbb{E}^{y}[\boldsymbol{t}|\Phi(\boldsymbol{x})]$ (see Section 3).
@@ D.1 Discrete Domain Case
For discrete domain problems, when there are a sufficient number of samples in each domain $\boldsymbol{t}$ and class $\boldsymbol{y}$, $\mathbb{E}^{y}[\boldsymbol{t}|\Phi(\boldsymbol{x})]$ and $\mathbb{E}^{t}[\boldsymbol{y}|\Phi(\boldsymbol{x})]$ performs similarly (both reflecting the first moment of $\boldsymbol{y}\perp\boldsymbol{t}|\Phi(\boldsymbol{x})$, although from different perspectives), and neither of them has a clear advantage over the other. Consider the following example, which has gained popularity [Lin et al., 2022b]: Consider a popular example [Lin et al., 2022b] as following: We have a binary classification task, $\boldsymbol{y}\in\{-1,1\}$, containing 2 domains $\boldsymbol{t}=\{1,2\}$, one invariant feature $\boldsymbol{x}_{v}$ and one spurious feature $\boldsymbol{x}_{s}$. Furthermore, we have
$\boldsymbol{x}_{v}=\begin{cases}1,\text{with prob }0.5,\\ -1,\text{with prob }0.5,\end{cases}\quad\boldsymbol{y}=\begin{cases}\boldsymbol{x}_{v},\text{ with prob }0.8,\\ -\boldsymbol{x}_{v},\text{ with prob }0.2,\end{cases}\text{ hold for both domains.}$. (3)
@@
In other words, the correlation between the invariant feature $\boldsymbol{x}_{v}$ and the label $\boldsymbol{y}$ is always 0.8. The correlation between $\boldsymbol{x}_{s}$ and $\boldsymbol{y}$ is 1 in domain 1 and 0.5 in domain 2. By simple calculation, we have
$\mathbb{E}^{t=1}[\boldsymbol{y}|\boldsymbol{x}_{s}=1]=1,\quad\mathbb{E}^{t=2}[\boldsymbol{y}|\boldsymbol{x}_{s}=1]=0$
$\mathbb{E}^{t=1}[\boldsymbol{y}|\boldsymbol{x}_{s}=-1]=-1,\quad\mathbb{E}^{t=2}[\boldsymbol{y}|\boldsymbol{x}_{s}=-1]=0$
$\mathbb{E}^{t=1}[\boldsymbol{y}|\boldsymbol{x}_{v}=1]=1,\quad\mathbb{E}^{t=2}[\boldsymbol{y}|\boldsymbol{x}_{v}=1]=1$
$\mathbb{E}^{t=1}[\boldsymbol{y}|\boldsymbol{x}_{v}=-1]=-1,\quad\mathbb{E}^{t=2}[\boldsymbol{y}|\boldsymbol{x}_{v}=-1]=-1$
@@
$\mathbb{E}^{y=1}[\boldsymbol{t}|\boldsymbol{x}_{s}=1]=4/3,\quad\mathbb{E}^{y=-1}[\boldsymbol{t}|\boldsymbol{x}_{s}=1]=2$
$\mathbb{E}^{y=1}[\boldsymbol{t}|\boldsymbol{x}_{s}=-1]=2,\quad\mathbb{E}^{y=-1}[\boldsymbol{t}|\boldsymbol{x}_{s}=-1]=4/3$
$\mathbb{E}^{y=1}[\boldsymbol{t}|\boldsymbol{x}_{v}=1]=3/2,\quad\mathbb{E}^{y=-1}[\boldsymbol{t}|\boldsymbol{x}_{v}=1]=3/2$
$\mathbb{E}^{y=1}[\boldsymbol{t}|\boldsymbol{x}_{v}=-1]=3/2,\quad\mathbb{E}^{y=-1}[\boldsymbol{t}|\boldsymbol{x}_{v}=-1]=3/2.$
So we can observe that $\mathbb{E}^{y}[\boldsymbol{t}|\boldsymbol{x}_{v}]$ is the same for all $\boldsymbol{y}$, and $\mathbb{E}^{t}[\boldsymbol{y}|\boldsymbol{x}_{v}]$ is the same for all $\boldsymbol{t}$. Additionally, $\mathbb{E}^{y}[\boldsymbol{t}|\boldsymbol{x}_{s}]$ is different for $y=-1$ and $y=1$, while $\mathbb{E}^{t}[\boldsymbol{y}|\boldsymbol{x}_{s}]$ is different for $t=1$ and $t=2$. In other words, if we aim to select a feature $\boldsymbol{x}$ from $\boldsymbol{x}\in\{\boldsymbol{x}_{v},\boldsymbol{x}_{s}\}$ to align either $\mathbb{E}^{y}[\boldsymbol{t}|\boldsymbol{x}]$ or $\mathbb{E}^{t}[\boldsymbol{y}|\boldsymbol{x}_{s}]$, we would choose the invariant feature $\boldsymbol{x}_{v}$ in both cases. So our CIL method and existing methods performs similarly in this example with 2 domains.
@@ D.2 Continuous Domain Case
Consider a 10-class classification task with an infinite number of domains and each domain contains only one sample, i.e., $n\xrightarrow[]{}\infty$, and $n/|\mathcal{T}|=1$, Proposition 1 shows that REx would fail to identify the invariant feature with constant probability. Whereas, Theorem 2 shows that CIL can still effectively extract invariant features. Intuitively, existing methods aim to align $\mathbb{E}^{t}[\boldsymbol{y}|\Phi(\boldsymbol{x})]$ across different $\boldsymbol{t}$ values. However, in continuous environment settings with limited samples per domain, the empirical estimations $\hat{\mathbb{E}}^{t}[\boldsymbol{y}|\Phi(\boldsymbol{x})]$ become noisy. These estimations deviate significantly from the true $\mathbb{E}^{t}[\boldsymbol{y}|\Phi(\boldsymbol{x})]$, rendering existing methods ineffective in identifying invariant features. In contrast, CIL proposes to align $\mathbb{E}^{y}[\boldsymbol{t}|\Phi(\boldsymbol{x})]$, which can be accurately estimated as there are sufficient sample in each class.
@@ Appendix F Proof of Proposition 1
If there are $n$ domains and each domain only contains one sample, then
@@
$=$ $\sum_{e}\mathcal{R}^{e}(\omega,\Phi)+\left(\sum_{e}\mathcal{R}^{e}(\omega,\Phi)-\sum_{e}\hat{\mathcal{R}}^{e}(\omega,\Phi)\right)$
$+\lambda\left(-\left(\mathcal{R}^{e}(\omega,\Phi)-\hat{\mathcal{R}}^{e}(\omega,\Phi)\right)+\left(\mathcal{R}(\omega,\Phi)-\hat{\mathcal{R}}(\omega,\Phi)\right)+\left(\mathcal{R}^{e}(\omega,\Phi)-\mathcal{R}(\omega,\Phi)\right)\right)^{2}$
$=$ $\sum_{e}\mathcal{R}^{e}(\omega,\Phi)+\lambda\underbrace{\sum_{e}\left(\mathcal{R}^{e}(\omega,\Phi)-\mathcal{R}(\omega,\Phi)\right)^{2}}_{A_{0}}$
$+\underbrace{\sum_{e}\left(\mathcal{R}^{e}(\omega,\Phi)-\hat{\mathcal{R}}^{e}(\omega,\Phi)\right)}_{A_{1}}+\underbrace{\sum_{e}\left(\mathcal{R}^{e}(\omega,\Phi)-\hat{\mathcal{R}}^{e}(\omega,\Phi)\right)^{2}}_{A_{2}}+\underbrace{\sum_{e}\left(\mathcal{R}(\omega,\Phi)-\hat{\mathcal{R}}(\omega,\Phi)\right)^{2}}_{A_{3}}$
$-\underbrace{\sum_{e}2\left(\mathcal{R}^{e}(\omega,\Phi)-\hat{\mathcal{R}}^{e}(\omega,\Phi)\right)\left(\mathcal{R}(\omega,\Phi)-\hat{\mathcal{R}}(\omega,\Phi)\right)}_{A_{4}}$
$-\underbrace{\sum_{e}2\left(\mathcal{R}^{e}(\omega,\Phi)-\hat{\mathcal{R}}^{e}(\omega,\Phi)\right)\left(\mathcal{R}^{e}(\omega,\Phi)-\mathcal{R}(\omega,\Phi)\right)}_{A_{5}}$
$+\underbrace{\sum_{e}2\left(\mathcal{R}(\omega,\Phi)-\hat{\mathcal{R}}(\omega,\Phi)\right)\left(\mathcal{R}^{e}(\omega,\Phi)-\mathcal{R}(\omega,\Phi)\right)}_{A_{6}}$
@@
$A_{0}=$$\begin{cases}\sigma_{R}\chi(|\mathcal{E}|),\mbox{ if }\Phi=\Phi_{s}\\ 0,\mbox{ if }\Phi=\Phi_{v}\\ 0,\mbox{ if }\Phi=\Phi_{null}\end{cases}$
$A_{1}=$ $\sum_{e}\sum_{i=1}^{n_{e}}\epsilon_{i}\sim\sigma/\sqrt{n}\mathcal{N}(0,1)$
$A_{2}=$ $\sum_{e}(\sum_{i=1}^{n_{e}}\epsilon_{i})^{2}\sim\sigma/\sqrt{n_{e}}\chi(|\mathcal{E}|)$
$A_{3}=$ $\sum_{e}(\sum_{i=1}^{n}\epsilon_{i})^{2}\sim\sigma/\sqrt{n}\chi(|\mathcal{E}|)$
$A_{4}=$ $2\left(\sum_{e}\left(\mathcal{R}^{e}(\omega,\Phi)-\hat{\mathcal{R}}^{e}(\omega,\Phi)\right)\right)\left(\mathcal{R}(\omega,\Phi)-\hat{\mathcal{R}}(\omega,\Phi)\right)$
@@
$A_{5}\sim$ $\sqrt{\sigma^{2}/\sqrt{n_{e}}+\sigma_{R}^{2}}\chi(|\mathcal{E}|)-\sqrt{\sigma^{2}/\sqrt{n_{e}}+\sigma_{R}^{2}}\chi(|\mathcal{E}|)$
$A_{6}=$ $0\displaystyle 2\left(\mathcal{R}(\omega,\Phi)-\hat{\mathcal{R}}(\omega,\Phi)\right)\sum_{e}\left(\mathcal{R}^{e}(\omega,\Phi)-\mathcal{R}(\omega,\Phi)\right)=2\left(\mathcal{R}(\omega,\Phi)-\hat{\mathcal{R}}(\omega,\Phi)\right)\times 0=$ 0
By taking account $n_{e}=n/|\mathcal{E}|$, we have
$\mathbb{E}[\hat{\mathcal{L}}(\Phi(\boldsymbol{x}))|\sigma_{s},\sigma_{v},\sigma_{null},\sigma_{R}]=\begin{cases}\sum_{e}\mathcal{R}^{e}(\omega,\Phi_{s})+\sigma_{R}|\mathcal{E}|+\frac{\sigma_{s}}{\sqrt{n}}(2+|\mathcal{E}|+|\mathcal{E}|^{2})\\ \sum_{e}\mathcal{R}^{e}(\omega,\Phi_{v})+\frac{\sigma_{v}}{\sqrt{n}}(2+|\mathcal{E}|+|\mathcal{E}|^{2})\\ \sum_{e}\mathcal{R}^{e}(\omega,\Phi_{null})+\frac{\sigma_{null}}{\sqrt{n}}(2+|\mathcal{E}|+|\mathcal{E}|^{2})\end{cases}$
Denote $Q=\left(\sum_{e}\mathcal{R}^{e}(w,\Phi_{v})-\sum_{e}\mathcal{R}^{e}(w,\Phi_{s})\right)$, assume $\delta_{R}|\mathcal{E}|\geq Q$. we have $\mathbb{E}[\hat{\mathcal{L}}(\Phi_{v}(\boldsymbol{x}))|\sigma_{s},\sigma_{v},\sigma_{R}]>\mathbb{E}[\hat{\mathcal{L}}(\Phi_{s}(\boldsymbol{x}))|\sigma_{s},\sigma_{v},\sigma_{R}]$ if
$(\sigma_{v}-\sigma_{s})\geq\frac{2\sigma_{R}\sqrt{n}}{|\mathcal{E}|}\geq\frac{\sqrt{n}(\sigma_{R}|\mathcal{E}|+|Q|)}{|\mathcal{E}|^{2}}\geq\frac{\sqrt{n}(\sigma_{R}|\mathcal{E}|+|Q|)}{(2+|\mathcal{E}|+|\mathcal{E}|^{2})}.$
@@
$|\mathcal{E}|\geq\frac{\sigma_{R}\sqrt{n}}{\Delta\mathcal{G}^{-1}(1/4)},$
with probability at least 1/4, REx can not identify the invariant feature.

### v2

@@ Appendix D Discussions on the first moments
To encourage conditional independence $\bm{y}\perp\bm{t}|\Phi(\bm{x})$, existing IRM variants try to align $\mathbb{E}^{t}[\bm{y}|\Phi(\bm{x})]$ (see Section 2.1) and our CIL proposes to align $\mathbb{E}^{y}[\bm{t}|\Phi(\bm{x})]$ (see Section 3).
@@ D.1 Discrete Domain Case
For discrete domain problems, when there are a sufficient number of samples in each domain $\bm{t}$ and class $\bm{y}$, $\mathbb{E}^{y}[\bm{t}|\Phi(\bm{x})]$ and $\mathbb{E}^{t}[\bm{y}|\Phi(\bm{x})]$ performs similarly (both reflecting the first moment of $\bm{y}\perp\bm{t}|\Phi(\bm{x})$, although from different perspectives), and neither of them has a clear advantage over the other. Consider the following example, which has gained popularity (Lin et al., 2022b): We have a binary classification task, $\bm{y}\in\{-1,1\}$, containing 2 domains $\bm{t}=\{1,2\}$, one invariant feature $\bm{x}_{v}$ and one spurious feature $\bm{x}_{s}$. Furthermore, we have
$\bm{x}_{v}=\begin{cases}1,\text{with prob }0.5,\\ -1,\text{with prob }0.5,\end{cases}\quad\bm{y}=\begin{cases}\bm{x}_{v},\text{ with prob }0.8,\\ -\bm{x}_{v},\text{ with prob }0.2,\end{cases}\text{ hold for both domains.}$. (3)
@@
In other words, the correlation between the invariant feature $\bm{x}_{v}$ and the label $\bm{y}$ is always 0.8. The correlation between $\bm{x}_{s}$ and $\bm{y}$ is 1 in domain 1 and 0.5 in domain 2. By simple calculation, we have
$\mathbb{E}^{t=1}[\bm{y}|\bm{x}_{s}=1]=1,\quad\mathbb{E}^{t=2}[\bm{y}|\bm{x}_{s}=1]=0$
$\mathbb{E}^{t=1}[\bm{y}|\bm{x}_{s}=-1]=-1,\quad\mathbb{E}^{t=2}[\bm{y}|\bm{x}_{s}=-1]=0$
$\mathbb{E}^{t=1}[\bm{y}|\bm{x}_{v}=1]=1,\quad\mathbb{E}^{t=2}[\bm{y}|\bm{x}_{v}=1]=1$
$\mathbb{E}^{t=1}[\bm{y}|\bm{x}_{v}=-1]=-1,\quad\mathbb{E}^{t=2}[\bm{y}|\bm{x}_{v}=-1]=-1$
@@
$\mathbb{E}^{y=1}[\bm{t}|\bm{x}_{s}=1]=4/3,\quad\mathbb{E}^{y=-1}[\bm{t}|\bm{x}_{s}=1]=2$
$\mathbb{E}^{y=1}[\bm{t}|\bm{x}_{s}=-1]=2,\quad\mathbb{E}^{y=-1}[\bm{t}|\bm{x}_{s}=-1]=4/3$
$\mathbb{E}^{y=1}[\bm{t}|\bm{x}_{v}=1]=3/2,\quad\mathbb{E}^{y=-1}[\bm{t}|\bm{x}_{v}=1]=3/2$
$\mathbb{E}^{y=1}[\bm{t}|\bm{x}_{v}=-1]=3/2,\quad\mathbb{E}^{y=-1}[\bm{t}|\bm{x}_{v}=-1]=3/2.$
So we can observe that $\mathbb{E}^{y}[\bm{t}|\bm{x}_{v}]$ is the same for all $\bm{y}$, and $\mathbb{E}^{t}[\bm{y}|\bm{x}_{v}]$ is the same for all $\bm{t}$. Additionally, $\mathbb{E}^{y}[\bm{t}|\bm{x}_{s}]$ is different for $y=-1$ and $y=1$, while $\mathbb{E}^{t}[\bm{y}|\bm{x}_{s}]$ is different for $t=1$ and $t=2$. In other words, if we aim to select a feature $\bm{x}$ from $\bm{x}\in\{\bm{x}_{v},\bm{x}_{s}\}$ to align either $\mathbb{E}^{y}[\bm{t}|\bm{x}]$ or $\mathbb{E}^{t}[\bm{y}|\bm{x}_{s}]$, we would choose the invariant feature $\bm{x}_{v}$ in both cases. So our CIL method and existing methods performs similarly in this example with 2 domains.
@@ D.2 Continuous Domain Case
Consider a 10-class classification task with an infinite number of domains and each domain contains only one sample, i.e., $n\xrightarrow[]{}\infty$, and $n/|\mathcal{T}|=1$, Proposition 1 shows that REx would fail to identify the invariant feature with constant probability. Whereas, Theorem 2 shows that CIL can still effectively extract invariant features. Intuitively, existing methods aim to align $\mathbb{E}^{t}[\bm{y}|\Phi(\bm{x})]$ across different $\bm{t}$ values. However, in continuous environment settings with limited samples per domain, the empirical estimations $\hat{\mathbb{E}}^{t}[\bm{y}|\Phi(\bm{x})]$ become noisy. These estimations deviate significantly from the true $\mathbb{E}^{t}[\bm{y}|\Phi(\bm{x})]$, rendering existing methods ineffective in identifying invariant features. In contrast, CIL proposes to align $\mathbb{E}^{y}[\bm{t}|\Phi(\bm{x})]$, which can be accurately estimated as there are sufficient sample in each class.
@@ Appendix F Proof of Proposition 1
Recall that $\mathcal{R}^{t}(\bm{w},\Phi):=\mathbb{E}^{t}[\ell(\bm{w}(\Phi(\bm{x})),\bm{y})]$ denotes the lose of $(\bm{w},\Phi)$ in domain $t$, where $\bm{w}$ is the classifier $\Phi$ is the feature selector, and $\ell$ is the lose function. If there are many domains (i.e., $\mathcal{T}$ is large) and each domain only contains limited sample (i.e., $n/|\mathcal{T}|$ is limited), then the empirical REx loss is as follows:
@@
$=$ $\sum_{t\in\mathcal{T}}\mathcal{R}^{t}(\omega,\Phi)+\left(\sum_{t\in\mathcal{T}}\mathcal{R}^{t}(\omega,\Phi)-\sum_{t\in\mathcal{T}}\hat{\mathcal{R}}^{t}(\omega,\Phi)\right)$
$+\lambda|\mathcal{T}|\left(-\left(\mathcal{R}^{t}(\omega,\Phi)-\hat{\mathcal{R}}^{t}(\omega,\Phi)\right)+\left(\mathcal{R}(\omega,\Phi)-\hat{\mathcal{R}}(\omega,\Phi)\right)+\left(\mathcal{R}^{t}(\omega,\Phi)-\mathcal{R}(\omega,\Phi)\right)\right)^{2}$
$=$ $\sum_{t\in\mathcal{T}}\mathcal{R}^{t}(\omega,\Phi)+\lambda\underbrace{\sum_{t\in\mathcal{T}}\left(\mathcal{R}^{t}(\omega,\Phi)-\mathcal{R}(\omega,\Phi)\right)^{2}}_{A_{0}}$
$+\underbrace{\sum_{t\in\mathcal{T}}\left(\mathcal{R}^{t}(\omega,\Phi)-\hat{\mathcal{R}}^{t}(\omega,\Phi)\right)}_{A_{1}}+\underbrace{\sum_{t\in\mathcal{T}}\left(\mathcal{R}^{t}(\omega,\Phi)-\hat{\mathcal{R}}^{t}(\omega,\Phi)\right)^{2}}_{A_{2}}+\underbrace{\sum_{t\in\mathcal{T}}\left(\mathcal{R}(\omega,\Phi)-\hat{\mathcal{R}}(\omega,\Phi)\right)^{2}}_{A_{3}}$
$-\underbrace{\sum_{t\in\mathcal{T}}2\left(\mathcal{R}^{t}(\omega,\Phi)-\hat{\mathcal{R}}^{t}(\omega,\Phi)\right)\left(\mathcal{R}(\omega,\Phi)-\hat{\mathcal{R}}(\omega,\Phi)\right)}_{A_{4}}$
$-\underbrace{\sum_{t\in\mathcal{T}}2\left(\mathcal{R}^{t}(\omega,\Phi)-\hat{\mathcal{R}}^{t}(\omega,\Phi)\right)\left(\mathcal{R}^{t}(\omega,\Phi)-\mathcal{R}(\omega,\Phi)\right)}_{A_{5}}$
$+\underbrace{\sum_{t\in\mathcal{T}}2\left(\mathcal{R}(\omega,\Phi)-\hat{\mathcal{R}}(\omega,\Phi)\right)\left(\mathcal{R}^{t}(\omega,\Phi)-\mathcal{R}(\omega,\Phi)\right)}_{A_{6}}$
@@
$A_{0}=$$\begin{cases}\sigma_{R}\chi(|\mathcal{T}|),\mbox{ if }\Phi=\Phi_{s}\\ 0,\mbox{ if }\Phi=\Phi_{v}\\ 0,\mbox{ if }\Phi=\Phi_{null}\end{cases}$
$A_{1}=$ $\sum_{t\in\mathcal{T}}\sum_{i=1}^{n_{e}}\epsilon_{i}\sim\sigma/\sqrt{n}\mathcal{N}(0,1)$
$A_{2}=$ $\sum_{t}(\sum_{i=1}^{n_{e}}\epsilon_{i})^{2}\sim\sigma/\sqrt{n_{e}}\chi(|\mathcal{T}|)$
$A_{3}=$ $\sum_{t}(\sum_{i=1}^{n}\epsilon_{i})^{2}\sim\sigma/\sqrt{n}\chi(|\mathcal{T}|)$
$A_{4}=$ $2\left(\sum_{t\in\mathcal{T}}\left(\mathcal{R}^{t}(\omega,\Phi)-\hat{\mathcal{R}}^{t}(\omega,\Phi)\right)\right)\left(\mathcal{R}(\omega,\Phi)-\hat{\mathcal{R}}(\omega,\Phi)\right)$
@@
$A_{5}\sim$ $\sqrt{\sigma^{2}/\sqrt{n_{e}}+\sigma_{R}^{2}}\chi(|\mathcal{T}|)-\sqrt{\sigma^{2}/\sqrt{n_{e}}+\sigma_{R}^{2}}\chi(|\mathcal{T}|)$
$A_{6}=$ $0\displaystyle 2\left(\mathcal{R}(\omega,\Phi)-\hat{\mathcal{R}}(\omega,\Phi)\right)\sum_{t\in\mathcal{T}}\left(\mathcal{R}^{t}(\omega,\Phi)-\mathcal{R}(\omega,\Phi)\right)=2\left(\mathcal{R}(\omega,\Phi)-\hat{\mathcal{R}}(\omega,\Phi)\right)\times 0=$ 0
By taking account $n_{e}=n/|\mathcal{T}|$, we have
$\mathbb{E}[\hat{\mathcal{L}}(\Phi(\bm{x}))|\sigma_{s},\sigma_{v},\sigma_{null},\sigma_{R}]=\begin{cases}\sum_{t\in\mathcal{T}}\mathcal{R}^{t}(\omega,\Phi_{s})+\sigma_{R}|\mathcal{T}|+\frac{\sigma_{s}}{\sqrt{n}}(2+|\mathcal{T}|+|\mathcal{T}|^{2})\\ \sum_{t\in\mathcal{T}}\mathcal{R}^{t}(\omega,\Phi_{v})+\frac{\sigma_{v}}{\sqrt{n}}(2+|\mathcal{T}|+|\mathcal{T}|^{2})\\ \sum_{t\in\mathcal{T}}\mathcal{R}^{t}(\omega,\Phi_{null})+\frac{\sigma_{null}}{\sqrt{n}}(2+|\mathcal{T}|+|\mathcal{T}|^{2})\end{cases}$
Denote $Q=\left(\sum_{t\in\mathcal{T}}\mathcal{R}^{t}(w,\Phi_{v})-\sum_{t\in\mathcal{T}}\mathcal{R}^{t}(w,\Phi_{s})\right)$, assume $\delta_{R}|\mathcal{T}|\geq Q$. we have $\mathbb{E}[\hat{\mathcal{L}}(\Phi_{v}(\bm{x}))|\sigma_{s},\sigma_{v},\sigma_{R}]>\mathbb{E}[\hat{\mathcal{L}}(\Phi_{s}(\bm{x}))|\sigma_{s},\sigma_{v},\sigma_{R}]$ if
$(\sigma_{v}-\sigma_{s})\geq\frac{2\sigma_{R}\sqrt{n}}{|\mathcal{T}|}\geq\frac{\sqrt{n}(\sigma_{R}|\mathcal{T}|+|Q|)}{|\mathcal{T}|^{2}}\geq\frac{\sqrt{n}(\sigma_{R}|\mathcal{T}|+|Q|)}{(2+|\mathcal{T}|+|\mathcal{T}|^{2})}.$
@@
$|\mathcal{T}|\geq\frac{\sigma_{R}\sqrt{n}}{\Delta\mathcal{G}^{-1}(1/4)},$
REx is unable to identify the invariant feature with a probability of at least 1/4.